OpenStudy (anonymous):

the number of gallons of water in a tank t minutes after the tank has started to drain is given by the function q(t) = 200(30-t)^2.

OpenStudy (anonymous):

find q'(10) and explain the significance of Q'(1) in the context of this situation.

OpenStudy (anonymous):

i found q'(t) = -1200+400t and q'(10 is 2800 galloons but im not sure what it represents.

OpenStudy (anonymous):

@Concentrationalizing please help

OpenStudy (anonymous):

q(t) = 200(30-t)^2 q'(t) = -400(30-t) q'(10) = -8000 When you take a derivative, you're finding a rate of change. So our deriative represents how the number of gallons of water changes over time. Saying q'(10) is saying that after 10 minutes, our volume of water has decreased by 8000 gallons. q'(1) is just saying that our volume has changed some amount after 1 minute

OpenStudy (anonymous):

@612skies